determine the condition for the roots of the equation ax2+bx+c=0 to differ by 2 ?

Grade:12th pass

1 Answers

Vijay Mukati
7 years ago
Dear Student,

Please follow the given steps to get your answer.

Let m and n are the two roots for the given quadratic eqn

Given, mod(m – n) = 2, m + n = -b/a and mn = c/a
Now using: (m-n)2 = (m+n)2 – 4mn and above values we can find the value of m and n.

Thanks, Vj